# Probabilistic Predictions of Process-Induced Deformation in Carbon/Epoxy Composites Using a Deep Operator Network

### v2

@@ 6 Uncertainty Quantification for DeepONet
UQ is crucial for assessing the reliability of DeepONet predictions, particularly with noisy or limited data. By explicitly characterizing confidence in the model outputs, UQ enables more robust and trustworthy deployment of operator-learning frameworks in practical applications [iglesias2013ensemble]. In this section, we present UQ results for DeepONet, focusing on epistemic uncertainty and EKI. We demonstrate how transfer learning under uncertainty can predict deformation histories, together with associated uncertainty bounds, for the experimental temperature profiles.
@@ 6.2 Ensemble Kalman Inversion
The ensemble Kalman filter (EnKF), originally introduced by Evensen in 1994 for data assimilation in time dependent dynamical systems, has become a widely used tool owing to its robustness, ease of implementation, and strong numerical performance [evensen1994sequential]. Building on this foundation, iterative ensemble Kalman methods have been developed for solving inverse problems. In this context, the inferred solution is constrained to the subspace spanned by the initial ensemble, which allows for error estimates relative to the best approximation achievable within that subspace. Numerical studies have shown that this ensemble-based, derivative free strategy can attain accuracy comparable to classical least squares approaches, and in many cases close to the optimal approximation.
EKI is a gradient free methodology specifically designed for inverse problems. Recent work [pensoneault2025uncertainty] applied EKI to the training of DeepONets and demonstrated that it achieves substantially faster convergence while providing more reliable uncertainty quantification than sampling based alternatives such as HMC. Motivated by these results, we integrate EKI into our operator learning framework to obtain efficient and informative uncertainty estimates for DeepONet.
@@ 7 Optimization of the cure temperature profile
Figure 13 illustrates the performance of DeepONet for optimizing the deformation under the constrained optimization problem (4). The predicted optimal cure schedule corresponds to the intermediate point $A=(t_{1},T_{1})\approx(1.61\pm 0.02\penalty 10000\ \text{min},\,133.01\pm 0.03\penalty 10000\ ^{\circ}\text{C})$. It is worth noting that this optimized case is very close to the optimal cure profile previously obtained in [18_Limaye2024].

### v3

@@ 6 Uncertainty Quantification for DeepONet
UQ is crucial for assessing the reliability of DeepONet predictions, particularly with noisy or limited data. By explicitly characterizing confidence in the model outputs, UQ enables more robust and trustworthy deployment of operator-learning frameworks in practical applications [18]. In this section, we present UQ results for DeepONet, focusing on epistemic uncertainty and EKI. We demonstrate how transfer learning under uncertainty can predict deformation histories, together with associated uncertainty bounds, for the experimental temperature profiles.
@@ 6.2 Ensemble Kalman Inversion
The ensemble Kalman filter (EnKF), originally introduced by Evensen in 1994 for data assimilation in time dependent dynamical systems, has become a widely used tool owing to its robustness, ease of implementation, and strong numerical performance [10]. Building on this foundation, iterative ensemble Kalman methods have been developed for solving inverse problems. In this context, the inferred solution is constrained to the subspace spanned by the initial ensemble, which allows for error estimates relative to the best approximation achievable within that subspace. Numerical studies have shown that this ensemble-based, derivative free strategy can attain accuracy comparable to classical least squares approaches, and in many cases close to the optimal approximation.
EKI is a gradient free methodology specifically designed for inverse problems. Recent work [47] applied EKI to the training of DeepONets and demonstrated that it achieves substantially faster convergence while providing more reliable uncertainty quantification than sampling based alternatives such as HMC. Motivated by these results, we integrate EKI into our operator learning framework to obtain efficient and informative uncertainty estimates for DeepONet.
@@ 7 Optimization of the cure temperature profile
Figure 13 illustrates the performance of DeepONet for optimizing the deformation under the constrained optimization problem (4). The predicted optimal cure schedule corresponds to the intermediate point $A=(t_{1},T_{1})\approx(1.61\pm 0.02\penalty 10000\ \text{min},\,133.01\pm 0.03\penalty 10000\ ^{\circ}\text{C})$. It is worth noting that this optimized case is very close to the optimal cure profile previously obtained in [32].